# Nearest Neighbor CCP-Based Molecular Sequence Analysis

### v1

@@ IV Experimental Setup
In this section, we describe the different datasets used for experiments. We also go through the baseline methods and evaluation metrics we use for the classification. A Windows 10 $64$-bit machine with an Intel(R) Core i5 processor operating at $2.10$ GHz and $32$ GB of memory is used for all experiments. Our pre-processed datasets and code are available online for reproducibility 111available in the published version.
We use three datasets in this study to assess the effectiveness of the proposed method. We employ t-distributed stochastic neighbor embedding (t-SNE) [29] to examine the data for any natural (hidden) grouping. The t-SNE plots for various embedding techniques are displayed in Figure 1,  2, and  3 for Protein Subcellular, Coronavirus Host, and Human DNA datasets, respectively. We can observe that t-SNE can group similar classes in the case of Autoencoder with CCP-NN.
@@
Spike2Vec Uses the sliding window (of size k) to get $k$-mers and its count in the sequence to generate feature vectors.  [18]
@@
String Kernel Kernel Matrix Designs $n\times n$ kernel matrix that can be used with kernel classifiers or with kernel PCA to get feature vector.  [19]
@@
SeqVec Pretrained Large Language Model (LLM) Takes biological sequences as input and fine-tunes the weights based on a pre-trained model to get final embedding.  [31]
@@ IV-A Baseline Methods
Among the baseline methods discussed in Table I, we selected $4$ popular embedding generation models, including One Hot Encoding (OHE) [16], Spike2Vec [18], PWM2Vec [17], and Autoencoder [30] to be used as input to both vanilla CCP and CCP-NN for dimensionality reduction. These methods were simple to use (in terms of implementation compared to complex models like SeqVec, TAPE, and Protein Bert), easy and fast to compute (compared to WDGRL, which is computationally expensive and takes long computational time), and generate embeddings directly, which can be used for dimensionality reduction (unlike String kernel, which generates a kernel matrix, which has to be converted to embeddings using kernel PCA, which could cause loss of information).
@@ IV-B Dataset Statistics
We use $3$ datasets for our experiments. The datasets employed are listed below:
Protein Subcellular Locations The dataset Protein Subcellular Locations Dataset we employ consists of $5959$ unaligned protein sequences, each corresponding to a different subcellular location [34]. The labels for the classification task are these subcellular sites and there are $11$ unique labels in our dataset. These labels correspond to the proteins of plant cells and fungal cells, while animal cells share all localizations with them. Table II provides the distribution of classes.
@@
- - Total 5959
Coronavirus Host The NIAD Virus Pathogen Database and Analysis Resource(ViPR) [35] and GISAID [36] are used to retrieve the Spike molecular sequences of CoVs for all of the hosts. Table III (in the supplementary material) comprises details about the $21$ host types with their counts of sequences that we gathered through the annotation of the total of $5558$ complete protein sequence.
@@ V-A Results For Protein Subcellular dataset
The classification results (averaged over $5$ runs) for the proposed CCP-NN and its comparison with the CCP approach for the Protein Subcellular dataset are shown in Table V. We can observe that the proposed CCP-NN outperforms the original CCP-based low-dimensional representation for all evaluation metrics and achieves a near-perfect predictive classification performance. The performance gain for CCP-NN (using OHE with Decision Tree classifier and using Autoencoder with Decision Tree classifier), compared to the best CCP-based results (Spike2Vec with Random Forest classifier) is $50.3\%$, which highlights a significant improvement in terms of predictive accuracy.
@@
TABLE V: Classification results (averaged over $5$ runs) for Protein Subcellular dataset using Nearest Neighbour CCP (CCP-NN) and CCP. The best value for each embedding is shown with the underline. The overall best value for each evaluation metric is shown in bold.
The comparison of the best performing proposed method from Table V (i.e. CCP-NN with Autoencoder) with the existing baseline models (without CCP or CPP-NN) is shown in Table VI. We can observe that the proposed method significantly outperforms all baselines for all evaluation metrics other than the training runtime. Specifically, in terms of average accuracy, the proposed method with Autoencoder embedding achieves $28\%$ improvement compared to the second best (i.e. Protein Bert, a pre-trained transformer-based model) and achieves a near-perfect average accuracy score in the case of the Protein Subcellular dataset.
@@
TABLE VI: Classification result comparisons (averaged over $5$ runs) for the best performing proposed method (i.e. CCP-NN with Autoencoder) with baselines on Protein Subcellular dataset. The best value for each embedding is shown underlined. The overall best value for each evaluation metric is shown in bold.
The standard deviation (SD) results (for $5$ runs) for the baselines and the proposed method are shown in Table VII for the Protein Subcellular dataset. We can observe that in the majority of the cases, the SD values are towards the lower end (i.e. $<0.02$), which shows that there is not much variation in the results for different experimental runs having a random train-test split.
@@ V-B Results For Coronavirus Host Data
The classification results (averaged over $5$ runs) for the proposed CCP-NN and its comparison with the CCP approach for the Coronavirus Host dataset are shown in Table VIII. The best values for each embedding method are underlined while overall best values among all methods are shown in bold. We can observe that the proposed CCP-NN outperforms the original CCP-based low-dimensional representation for all evaluation metrics (other than classifier training runtime) and achieves a near-perfect predictive classification performance. Although the performance gain for CCP-NN compared to the best CCP-based results is not significant, however, it still outperforms CCP for all evaluation metrics other than training runtime.
The comparison of the best performing proposed method from Table VIII i.e. CCP-NN with Spike2Vec, with the existing baseline models is shown in Table IX for the Coronavirus Host dataset. We can observe that the proposed method significantly outperforms all baselines for all evaluation metrics other than the training runtime. Specifically, in terms of average accuracy, the CCP-NN with Spike2Vec embedding achieves $1.7\%$ improvement compared to the second-best results (i.e. original Spike2Vec with Random Forest and Logistic Regression classifiers) and achieves a higher average accuracy score in the case of the Coronavirus Host dataset.
@@
TABLE VIII: Classification results (averaged over $5$ runs) on Coronavirus Host dataset for different evaluation metrics using Nearest Neighbour CCP (CCP-NN) and CCP. The best value for each embedding is shown with the underline. The overall best value for each evaluation metric is shown in bold.
@@
TABLE IX: Classification result comparisons (averaged over $5$ runs) for the best performing proposed method (i.e., CCP-NN with Spike2Vec) with baselines on Coronavirus Host dataset for different evaluation metrics. The best values are in bold.
Table X displays the standard deviation (SD) outcomes (averaged over 5 runs) for both the baseline methods and proposed approach on the Coronavirus Host dataset. The results indicate that, in most cases, the SD values are relatively low, usually below 0.02. This observation suggests that the classification results remain consistent across different experimental runs with random train-test splits. The low variability in the SD values reflects the stability of the reported classification performance for both the proposed and baseline models.
@@ V-C Results For Human DNA Data
TABLE XI: Classification results (averaged over $5$ runs) on Human DNA dataset for different evaluation metrics using Nearest Neighbour CCP (CCP-NN) and CCP. The best value for each embedding is shown with the underline. The overall best value for each evaluation metric is shown in bold.
@@
TABLE XII: Classification result comparisons (averaged over $5$ runs) for the best performing proposed method (i.e. CCP-NN with Autoencoder) with baselines on Human DNA dataset for different evaluation metrics. The best values are in bold.
The comparison of the best performing proposed method from Table XI, i.e. CCP-NN with Autoencoder, with the existing baseline models is shown in Table XII for the Human DNA dataset. We can observe that the proposed method significantly outperforms all baselines for all evaluation metrics other than the training runtime. Specifically, in terms of average accuracy, the CCP-NN with Autoencoder embedding achieves $11.8\%$ improvement compared to the second-best results (i.e. original Spike2Vec with Random Forest classifier).
@@ V-D Runtime Evaluation
[…]
@@
The computational runtime for $R_{\phi_{CCP\_NN}}$ and $R_{\phi_{CCP}}$ along with the performance gain is reported in Table XVI,  XVI, and  XVI for the Protein Subcellular, Human DNA, and Coronavirus Host datasets, respectively. For protein subcellular data, we can observe that for all $4$ embedding methods as input to the CCP and CCP-NN, the performance gain (i.e. Percentage improvement) for our CCP-NN is $23.32\%$, $72.34\%$, $55.93\%$, and $92.88\%$, for OHE, Spike2Vec, PWM2Vec, and Autoencoder, respectively. For the Coronavirus Host dataset, we can again observe that in terms of computations runtime performance gain, the proposed CCP-NN significantly outperforms the original CCP by $39.429\%$, $55.950\%$, $97.865\%$, and $93.989\%$ for OHE, Spike2Vec, PWM2Vec, and Autoencoder, respectively. Similarly, for the Human DNA dataset, we can observe $94.385\%$ (OHE), $90.243\%$ (for Spike2Vec), $91.456\%$ (PWM2Vec), and $85.425\%$ (for Autoencoder) improvement in computational runtime for CCP-NN compared to the CCP.
To observe the increase in runtime with the increasing number of data points (i.e. embeddings), we select the overall best-performing embedding method with CCP and CCP-NN, i.e. Autoencoder, and compute runtime with an increasing number of embeddings. The runtime results are reported in Figure 4. We can observe that the vanilla CCP’s runtime increases very quickly as we increase the number of embeddings. On the other hand, the runtime increase for the CCP-NN is very slow, showing its better scalability property.
Embedding $\phi_{CCP\_NN}$ (ours) (in seconds) $\phi_{CPP}$ (in seconds) % Improvement for $\phi_{CCP\_NN}$ over $\phi_{CPP}$
@@
Embedding $\phi_{CCP\_NN}$ (ours) (in seconds) $\phi_{CPP}$ (in seconds) % Improvement for $\phi_{CCP\_NN}$ over $\phi_{CPP}$
@@
Embedding $\phi_{CCP\_NN}$ (ours) (in seconds) $\phi_{CPP}$ (in seconds) % Improvement for $\phi_{CCP\_NN}$ over $\phi_{CPP}$
@@ V-E Statistical Significance
We used the Student’s t-test and calculated the $p$-values using the average and standard deviations (SD) of $5$ runs to determine whether the computed classification results are statistically significant. Given that SD values are extremely low for all datasets, i.e. $<0.002$ in most cases, we noticed that $p$-values were $<0.05$ in the majority of cases and for all embedding methods, hence validating the statistical significance of the findings.
@@ V-F Discussion
The proposed method employs the nearest neighbor (NN) search technique using the AnnoyIndex. By leveraging NN, the algorithm reduces the computational complexity of pairwise distance calculations. In high-dimensional spaces like biological sequencing data where feature dimensions can be large, the NN-based approach significantly speeds up the computation, allowing for better handling of the complexity and potentially better capturing of correlations between features.
@@
In sequence classification, feature spaces can often be high-dimensional due to the representation of various attributes. The proposed method can better preserve the feature information in the low-dimensional space than the original CCP due to efficient nearest-neighbor computation, leading to more meaningful density estimations. The nearest neighbor search focuses on capturing local patterns in data, which are particularly relevant for correlations between features. Using the density estimation from an NN-based neighbor search, the method effectively identifies relationships between relevant features, thereby representing the underlying patterns better.
@@
The data might contain noise and outliers, which can negatively impact the accuracy of correlation estimation. The NN-based method can be more robust to noisy data as it focuses on local patterns rather than global distances. It implicitly handles noise using neighbor relationships, leading to more reliable density estimations and better classification results.
@@
The proposed method partitions the features into clusters based on their correlation patterns. By employing NN-based density estimation, the method identifies more optimal feature clusters, which in turn can enhance the separation of target classes. The ability to detect relevant feature subsets for classification can contribute to improved accuracy.
In general, the superior classification results obtained with the proposed method can be attributed to its efficient handling of high-dimensional data, the use of NN for nearest neighbor search, enhanced correlation estimation, robustness to noise, and optimal feature clustering. These advantages collectively enable the method to capture the underlying patterns and provide more discriminative representations for improved classification performance as demonstrated in a variety of datasets including protein subcellular localization data, Coronavirus Host data, and Human DNA in our experiments.

### v2

@@ IV Experimental Setup
In this section, we describe the different datasets used for the experiments. We also go through the baseline methods and evaluation metrics we use for the classification. A Windows 10 $64$-bit machine with an Intel(R) Core i5 processor operating at $2.10$ GHz and $32$ GB of memory is used for all experiments. We used standard cross-validation to tune the hyperparameters.
In this study, we used three datasets to assess the effectiveness of the proposed method. We employ t-distributed stochastic neighbor embedding (t-SNE) [37] to examine the data for any natural (hidden) grouping. The t-SNE plots for various embedding techniques are shown in Figure 1,  2, and  3 for Protein Subcellular, Coronavirus Host, and Human DNA datasets, respectively. We can observe that t-SNE can group similar classes in the case of the Autoencoder with CCP-NN.
@@
Spike2Vec Uses the sliding window (of size k) to get $k$-mers and their count in the sequence to generate feature vectors.  [25]
@@
String Kernel Kernel Matrix Designs $n\times n$ kernel matrix that can be used with kernel classifiers or with kernel PCA to get a feature vector.  [26]
@@
SeqVec Pretrained Large Language Model (LLM) Takes biological sequences as input and fine-tunes the weights based on a pre-trained model to get the final embedding.  [39]
@@ IV-A Baseline Methods
Among the baseline methods discussed in Table I, we selected $4$ popular embedding generation models, including One Hot Encoding (OHE) [21], Spike2Vec [25], PWM2Vec [22], and Autoencoder [38] to be used as input to both vanilla CCP and CCP-NN for dimensionality reduction. These methods were simple to use (in terms of implementation compared to complex models like SeqVec, TAPE, and Protein Bert), easy and fast to compute (compared to WDGRL, which is computationally expensive and takes a long computational time), and generate embeddings directly, which can be used for dimensionality reduction (unlike String kernel, which generates a kernel matrix, which has to be converted to embeddings using kernel PCA, which could cause loss of information).
@@ IV-B Dataset Statistics
We use $3$ datasets for our experiments. The effectiveness of the k-NN algorithm is heavily dependent on the quality and scale of the dataset, as it uses distance-based calculations. To address this, we performed data normalization during preprocessing to ensure that all features contribute equally, improving the accuracy and reliability of the model. The datasets used are listed below:
Protein Subcellular Locations The Protein Subcellular Locations dataset we used consists of $5959$ unaligned protein sequences, each corresponding to a different subcellular location [42]. The labels for the classification task are these subcellular sites, and there are $11$ unique labels in our dataset. These labels correspond to the proteins of plant cells and fungal cells, while animal cells share all localizations with them. Table II provides the distribution of classes.
@@
- - Total $59$
Coronavirus Host The NIAD Virus Pathogen Database and Analysis Resource(ViPR) [43] and GISAID [44] are used to retrieve Spike molecular sequences of CoVs for all hosts. Table III (in the supplementary material) comprises details about the $21$ host types with their counts of sequences that we gathered through the annotation of the total of $5558$ complete unaligned spike protein sequences. We have some hosts that have very few sequences; these sequences ensure that the dataset represents the full diversity of known hosts rather than omitting certain species entirely. In addition, a single sequence may still carry host-specific signatures that could be useful for feature analysis.
TABLE III: Statistics for Coronavirus Host dataset.
@@ V-A Results For Protein Subcellular dataset
The classification results (averaged over $5$ runs) for the proposed CCP-NN and its comparison with the CCP approach for the Protein Subcellular dataset are shown in Table V. We can observe that the proposed CCP-NN outperforms the original CCP-based low-dimensional representation for all evaluation metrics and achieves a near-perfect predictive classification performance. The performance gain for CCP-NN (using OHE with the Decision Tree classifier and using Autoencoder with the Decision Tree classifier), compared to the best CCP-based results (Spike2Vec with the Random Forest classifier), is $50.3\%$, highlighting a significant improvement in terms of predictive accuracy.
TABLE V: Classification results (averaged over $5$ runs) for Protein Subcellular dataset using Nearest Neighbour CCP (CCP-NN) and CCP. The best value for each embedding is shown with an underline. The overall best value for each evaluation metric is shown in bold. The $\uparrow$ means a higher number is better, and the down arrow $\downarrow$ means that a lower number is better.
@@
The comparison of the best performing proposed method from Table V (i.e., CCP-NN with Autoencoder) with the existing baseline models (without CCP or CPP-NN) is shown in Table VI. We can observe that the proposed method significantly outperforms all baselines for all evaluation metrics other than the training runtime. Specifically, in terms of average accuracy, the proposed method with Autoencoder embedding achieves $28\%$ improvement compared to the second best (i.e., Protein Bert, a pre-trained transformer-based model) and achieves a near-perfect average accuracy score in the case of the Protein Subcellular dataset.
TABLE VI: Classification result comparisons (averaged over $5$ runs) for the best performing proposed method (i.e., CCP-NN with Autoencoder) with baselines on Protein Subcellular dataset. The best value for each embedding is shown in underlined. The overall best value for each evaluation metric is shown in bold. The $\uparrow$ means a higher number is better, and the down arrow $\downarrow$ means that a lower number is better.
@@
The standard deviation (SD) results (for $5$ runs) for the baselines and the proposed method are shown in Table VII for the Protein Subcellular dataset. We can observe that in the majority of the cases, the SD values are towards the lower end (i.e. <0.02<0.02), which shows that there is not much variation in the results for different experimental runs having a random train-test split.
@@ V-B Results For Coronavirus Host Data
The classification results (averaged over $5$ runs) for the proposed CCP-NN and its comparison with the CCP approach for the Coronavirus Host dataset are shown in Table VIII. The best values for each embedding method are underlined, while the overall best values among all methods are shown in bold. We can observe that the proposed CCP-NN outperforms the original CCP-based low-dimensional representation for all evaluation metrics (other than the classifier training runtime) and achieves near-perfect predictive classification performance. Although the performance gain for CCP-NN compared to the best CCP-based results is not significant, it still outperforms CCP for all evaluation metrics other than training runtime.
The comparison of the best performing proposed method from Table VIII, i.e., CCP-NN with Spike2Vec, with the existing baseline models is shown in Table IX for the Coronavirus Host dataset. We can observe that the proposed method significantly outperforms all baselines for all evaluation metrics other than the training runtime. Specifically, in terms of average accuracy, the CCP-NN with Spike2Vec embedding achieves a $1.7\%$ improvement compared to the second best results (i.e., original Spike2Vec with Random Forest and Logistic Regression classifiers) and achieves a higher average accuracy score in the case of the Coronavirus Host dataset.
TABLE VIII: Classification results (averaged over $5$ runs) on Coronavirus Host dataset for different evaluation metrics using Nearest Neighbour CCP (CCP-NN) and CCP. The best value for each embedding is shown with an underline. The overall best value for each evaluation metric is shown in bold. The $\uparrow$ means a higher number is better, and the down arrow $\downarrow$ means that a lower number is better.
@@
TABLE IX: Classification result comparisons (averaged over $5$ runs) for the best performing proposed method (i.e., CCP-NN with Spike2Vec) with baselines on Coronavirus Host dataset for different evaluation metrics. The best values are in bold. The $\uparrow$ means a higher number is better, and the down arrow $\downarrow$ means that a lower number is better.
@@
Table X displays the results of the standard deviation (SD) (averaged over 5 runs) for both the baseline methods and the approach proposed in the Coronavirus Host dataset. The results indicate that in most cases the SD values are relatively low, usually below 0.02. This observation suggests that the classification results remain consistent between different experimental runs with random train-test splits. The low variability in the SD values reflects the stability of the classification performance reported for both the proposed and the baseline models.
@@ V-C Results For Human DNA Data
TABLE XI: Classification results (averaged over $5$ runs) on Human DNA dataset for different evaluation metrics using Nearest Neighbour CCP (CCP-NN) and CCP. The best value for each embedding is shown with an underline. The overall best value for each evaluation metric is shown in bold. The $\uparrow$ means a higher number is better, and the down arrow $\downarrow$ means that a lower number is better.
@@
TABLE XII: Classification result comparisons (averaged over $5$ runs) for the best performing proposed method (i.e., CCP-NN with Autoencoder) with baselines on Human DNA dataset for different evaluation metrics. The best values are in bold. The $\uparrow$ means a higher number is better, and the down arrow $\downarrow$ means that a lower number is better.
@@
The comparison of the best performing proposed method from Table XI, i.e., CCP-NN with Autoencoder, with the existing baseline models is shown in Table XII for the Human DNA dataset. We can observe that the proposed method significantly outperforms all baselines for all evaluation metrics other than the training runtime. Specifically, in terms of average accuracy, the CCP-NN with Autoencoder embedding achieves $11.8\%$ improvement compared to the second-best results (i.e., original Spike2Vec with Random Forest classifier).
@@ V-D Runtime Evaluation
For all datasets, we additionally report $\%$ improvement for running the $\phi_{CCP\_NN}$ compared to $\phi_{CCP}$ in terms of runtime. For computing the runtime performance gain, we use the following expression:
% $\text{\% improvement}\;\;=\;\;\dfrac{\text{$R_{\phi_{CCP}}$ $-$ $R_{\phi_{CCP\_NN}}$}}{\text{$R_{\phi_{CCP}}$}}\times 100$ (15)
@@
The computational runtime for $R_{\phi_{CCP\_NN}}$ and $R_{\phi_{CCP}}$ along with the performance gain is reported in Tables XIV,  XV, and  XVI for the Protein Subcellular, Human DNA and Coronavirus Host datasets, respectively. For protein subcellular data, we can observe that for all $4$ embedding methods as input to CCP and CCP-NN, the performance gain (i.e., percentage improvement) for our CCP-NN is $23.32\%$, $72.34\%$, $55.93\%$, and $92.88\%$, for OHE, Spike2Vec, PWM2Vec, and Autoencoder, respectively. For the Coronavirus Host dataset, We can again observe that in terms of computation runtime performance gain, the proposed CCP-NN significantly outperforms the original CCP by $39.429\%$, $55.950\%$, $97.865\%$, and $93.989\%$ for OHE, Spike2Vec, PWM2Vec, and Autoencoder, respectively. Similarly, for the Human DNA dataset, We can observe a $94.385\%$ (OHE), $90.243\%$ (for Spike2Vec), $91.456\%$ (PWM2Vec), and $85.425\%$ (for Autoencoder) improvement in computational runtime for CCP-NN compared to CCP.
To observe the increase in runtime with the increasing number of data points (i.e., embeddings), we select the overall best-performing embedding method with CCP and CCP-NN, i.e., Autoencoder, and compute the runtime with an increasing number of embeddings. The runtime results are reported in Figure 4. We can observe that the vanilla CCP’s runtime increases very quickly as we increase the number of embeddings. On the other hand, the runtime increase for the CCP-NN is very slow, showing its better scalability property.
Embedding $\phi_{CCP\_NN}$
in seconds
$\phi_{CPP}$
in seconds
% Improvement for $\phi_{CCP\_NN}$ over $\phi_{CPP}$
@@
Embedding $\phi_{CCP\_NN}$
in seconds
$\phi_{CPP}$
in seconds
% Improvement for $\phi_{CCP\_NN}$ over $\phi_{CPP}$
@@
Embedding $\phi_{CCP\_NN}$
in seconds
$\phi_{CPP}$
in seconds
% Improvement for $\phi_{CCP\_NN}$ over $\phi_{CPP}$
@@ V-E Statistical Significance
We used the Student’s t-test and calculated the $p$-values using the average and standard deviations (SD) of $5$ runs to determine whether the computed classification results are statistically significant. Given that SD values are extremely low for all datasets, i.e., <0.002<0.002 in most cases, we noticed that $p$-values were <0.05<0.05 in the majority of cases and for all embedding methods, hence validating the statistical significance of the findings.
@@ V-F Discussion
The proposed method employs the nearest neighbor (NN) search technique using the AnnoyIndex. Using NN, the algorithm reduces the computational complexity of pairwise distance calculations. In high-dimensional spaces such as biological sequencing data, where feature dimensions can be large, the NN-based approach significantly speeds up the computation, allowing better handling of the complexity and potentially better capturing of correlations between features.
@@
In sequence classification, feature spaces can often be high-dimensional due to the representation of various attributes. The proposed method can better preserve the feature information in the low-dimensional space than the original CCP due to efficient nearest-neighbor computation, leading to more meaningful density estimations. The nearest neighbor search focuses on capturing local patterns in the data that are particularly relevant for correlations between features. Using the density estimation from an NN-based neighbor search, the method effectively identifies the relationships between relevant features, thus better representing the underlying patterns.
@@
Data may contain noise and outliers, which can negatively impact the accuracy of the correlation estimation. The NN-based method can be more robust to noisy data, as it focuses on local patterns rather than global distances. It implicitly handles noise using neighbor relationships, leading to more reliable density estimates and better classification results.
@@
The proposed method partitions the features into clusters based on their correlation patterns. Using NN-based density estimation, the method identifies more optimal feature clusters, which in turn can enhance the separation of target classes. The ability to detect relevant feature subsets for classification can contribute to improved accuracy.
In general, the superior classification results obtained with the proposed method can be attributed to its efficient handling of high-dimensional data, the use of NN for nearest neighbor search, enhanced correlation estimation, robustness to noise, and optimal feature clustering. These advantages collectively enable the method to capture the underlying patterns and provide more discriminative representations for improved classification performance, as demonstrated in a variety of datasets, including protein subcellular localization data, Coronavirus host data, and Human DNA in our experiments.
V-G Results Comparisons on Gene Expression Datasets
To evaluate the performance of the proposed CCP_NNCCP\_NN method compared to the original CCP method without relying on any specific type of embeddings, we conducted experiments using two gene expression datasets: Leukemia [46] and ALL-AML [47]. The Leukemia dataset includes microarray gene expression values from leukemia $72$ patients, where each patient has $7070$ gene expression values. The binary classification task differentiates between “myeloid” and “lymphoblastic” leukemia. The ALL-AML dataset consists of $7129$ gene expression profiles of $72$ patients with acute lymphoblastic leukemia (ALL) and acute myeloid leukemia (AML), hence the binary cancer classification task is performed.
The results are summarized in Tables XVII and XVIII for the Leukemia and ALL-AML datasets, respectively. We can observe that the CCP_NNCCP\_NN method demonstrates competitive performance compared to the original CCP method for both datasets. This behavior shows that even without relying on a specific embedding generation method, the proposed method exhibits substantial improvements over the original CCP method, demonstrating its potential in gene expression analysis and classification tasks. The detailed comparisons and results are provided in the tables, where the best values for each method are underlined, and the overall best values are highlighted in bold.
TABLE XVII: Classification result comparisons (averaged over $5$ runs) for $\phi_{CCP}$ and $\phi_{CCP\_NN}$ on Lukemia Gene Expression dataset for different evaluation metrics. The best values for each method are underlined, while the overall best values are in bold. The $\uparrow$ means a higher number is better, and the down arrow $\downarrow$ means that a lower number is better.
$\phi_{CCP}$ SVM 0.909 0.923 0.909 0.910 0.902 0.911 0.002
NB 0.900 0.926 0.900 0.901 0.896 0.910 0.002
MLP 0.882 0.907 0.882 0.882 0.874 0.889 0.117
KNN 0.864 0.879 0.864 0.865 0.855 0.866 0.003
RF 0.900 0.925 0.900 0.901 0.895 0.912 0.188
LR 0.900 0.914 0.900 0.901 0.892 0.899 0.006
DT 0.745 0.781 0.745 0.731 0.711 0.718 0.015
$\phi_{CCP\_NN}$ (ours) SVM 0.691 0.508 0.691 0.582 0.408 0.489 0.003
NB 0.618 0.728 0.618 0.629 0.596 0.677 0.001
MLP 0.809 0.851 0.809 0.798 0.737 0.738 0.152
KNN 0.827 0.844 0.827 0.823 0.772 0.770 0.003
RF 0.936 0.947 0.936 0.938 0.919 0.944 0.210
LR 0.691 0.508 0.691 0.582 0.408 0.489 0.005
DT 0.945 0.954 0.945 0.947 0.926 0.948 0.017
TABLE XVIII: Classification result comparisons (averaged over $5$ runs) for $\phi_{CCP}$ and $\phi_{CCP\_NN}$ on ALL-AML dataset for different evaluation metrics. The best values for each method are underlined, while the overall best values are in bold. The $\uparrow$ means a higher number is better, and the down arrow $\downarrow$ means that a lower number is better.
$\phi_{CCP}$ SVM 0.955 0.959 0.955 0.954 0.952 0.950 0.004
NB 0.936 0.940 0.936 0.936 0.931 0.932 0.003
MLP 0.936 0.937 0.936 0.936 0.934 0.936 0.204
KNN 0.945 0.952 0.945 0.944 0.938 0.928 0.005
RF 0.945 0.952 0.945 0.944 0.940 0.934 0.487
LR 0.964 0.967 0.964 0.963 0.959 0.951 0.009
DT 0.900 0.912 0.900 0.900 0.891 0.893 0.039
$\phi_{CCP\_NN}$ (ours) SVM 0.727 0.732 0.727 0.727 0.676 0.677 0.004
NB 0.673 0.690 0.673 0.678 0.628 0.636 0.002
MLP 0.855 0.867 0.855 0.854 0.818 0.822 0.564
KNN 0.964 0.971 0.964 0.965 0.955 0.969 0.005
RF 0.955 0.958 0.955 0.955 0.946 0.951 0.465
LR 0.691 0.483 0.691 0.567 0.407 0.500 0.009
DT 0.945 0.950 0.945 0.945 0.932 0.935 0.045